3+√(z-5)=√(z+10)
please help me solve this equation thanks

Answer by stanbon(75887)

3+√(z-5)=√(z+10)
---
Square both sides to get:
9 + 6sqrt(z-5) + (z-5) = z+10
---
Rearrange:
6sqrt(z-5) = 6
----
sqrt(z-5) = 1
Square both sides:
z-5 = 1
z = 6
-----
Check the "answer":
3 + sqrt(6-5) = sqrt(6+10)
3 + 1 = 4
4 = 4
---------------
Therefore: z = 6
=======================
Cheers,
Stan H.